OpenStudy (anonymous):

a plane flies 480 miles with wind 320 mile against wind the same length of time if the wind speed 22mph whats the speed of the plane in still air.

OpenStudy (sgadi):

lets take x as the speed of flight and t as the time. (x+22)*t=480 (x-22)*t=320 by solving above equations , we can get t and x \[ t=\frac{40}{11} \] \[x=110\]

OpenStudy (anonymous):

did you plug in x with a number than subtract

OpenStudy (sgadi):

yes....